OpenStudy (anonymous):

Suppose that the position of a particle as a function of time (in seconds)is given by the formula s(t)=5.5+5t^3-t^5, t>0 a)The time at which the velocity is zero is b)The time at which the acceleration is zero is c)The time at which the velocity is maximum is

OpenStudy (anonymous):

velocity is ds/dt,time at which the velocity is zero is the time at which ds/dt=0

OpenStudy (anonymous):

can you show me how to do it, im confused

OpenStudy (anonymous):

ds/dt=5.5+15t^2-5t^4 is that understood?

OpenStudy (anonymous):

sorry ds/dt=15t^2-5t^4

OpenStudy (anonymous):

but you need to understand why

OpenStudy (anonymous):

yea its the derrivitve

OpenStudy (anonymous):

okay ..then we need 15t^2-5t^4=0

OpenStudy (anonymous):

so itll be 5t^2(3-t^2)=0

OpenStudy (anonymous):

or better 5(3t^2-t^4)=0

OpenStudy (anonymous):

and then because 5 cant be zero we need 3t^2-t^4 to be zero

OpenStudy (anonymous):

so 3t^2-t^4=0 ----->3t^2=t^4 ok till now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3t^2=t^4 ----->3=t^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[t=\sqrt{3}\] or \[t=-\sqrt{3}\]

OpenStudy (anonymous):

ok till now?

OpenStudy (anonymous):

yese

OpenStudy (anonymous):

can we have negative time? :o

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so its sqrt(3)

OpenStudy (anonymous):

then we have \[t=\sqrt{3}\] seconds

OpenStudy (anonymous):

thanks... for the acc. do we set the second derr. to zero?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

30t-20t^3=0 10(3-2t^2)=0

OpenStudy (anonymous):

3-2t^2=0 3=2t^2 ans sqrt(3/2)

OpenStudy (anonymous):

how do we find part c?

OpenStudy (anonymous):

good job on part 2

OpenStudy (anonymous):

how do we find part c?

OpenStudy (anonymous):

umax=at right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the maximum speed is acceleration*time

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so what do we do next

OpenStudy (anonymous):

accelaration*time=30-20t^3*t and velocity=15t^2-5t^4

OpenStudy (anonymous):

sorry acceleration*time=(30-20t^3)*t

OpenStudy (anonymous):

damn (30t-20t^3)*t

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

30t^2-20t^4=15t^2-5t^4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then we divide right and left hand side with t^2

OpenStudy (anonymous):

Hmmm? Use the velocity equation to find the velocity at a particular time.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Use acceleration equation to find critical points in time.

OpenStudy (anonymous):

max and min are at critical points.

OpenStudy (anonymous):

yes but because u=at we have u and a so dont we find t from there?

OpenStudy (anonymous):

The purpose of b) was to find \(t\).

OpenStudy (anonymous):

oh damn yeah my bad i see what you mean

OpenStudy (anonymous):

It's an open interval so there is only one critical point and if that isn't the maximum then there is no maximum. Open intervals are not subject to extreme value theorem.

OpenStudy (anonymous):

so what do we do?

OpenStudy (anonymous):

Plug in \(t\) from b) into your velocity equation.

OpenStudy (anonymous):

you need to put sqrt(3/2) into 5(3t^2-t^4)

OpenStudy (anonymous):

and then set it equalto zero ?

OpenStudy (anonymous):

no you ll find a number..that's the maximum velocity

OpenStudy (anonymous):

wait im confused

OpenStudy (anonymous):

we multpile them?

OpenStudy (anonymous):

you put where t is at the velocity equation the acceleration we found at question b that is sqrt(3/2)

OpenStudy (anonymous):

so maximum speed=15*sqrt(3/2)^2-5sqrt(3/2)^4

OpenStudy (anonymous):

you put the result of our b question=acceleration into our velocity equation we got from question a

OpenStudy (anonymous):

what answer are you getting as the final, its saying my answer is wrong

OpenStudy (anonymous):

yes i see why

OpenStudy (anonymous):

i got a bit confused with question c...brb checking my old kinematic notes

OpenStudy (anonymous):

k

OpenStudy (anonymous):

still seems the procedure is right...what result did u find?

OpenStudy (anonymous):

i just entered it exactly like that before the final answer and it says inccorect

OpenStudy (anonymous):

what are you getting?

OpenStudy (anonymous):

the final answer is 11.43?

OpenStudy (anonymous):

no its coming up wrong

OpenStudy (anonymous):

part b is correct im just having a problem with part c

OpenStudy (anonymous):

yes it is i dont know i still get 11.43

OpenStudy (anonymous):

whats the whole number you got... any numbers after .43?

OpenStudy (anonymous):

basically its 11,271

OpenStudy (anonymous):

i redid it

OpenStudy (anonymous):

so would be 11.3

OpenStudy (anonymous):

its still coming up wrong.. ill just leave that for now

OpenStudy (anonymous):

thanks so much though!

OpenStudy (anonymous):

how much is the answer?

OpenStudy (anonymous):

ah wait

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i dont know the answer yet, did you find it?

OpenStudy (anonymous):

we found the max velocity,not the time

OpenStudy (anonymous):

(30t-20t^3)t=11.3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

30t^2-20t^4=11.3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i find imaginary parts in result ,must be wrong,sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

seems like the velocity is at a maximum the time acceleration is zero that means t=3/2=1.5 sec